OpenStudy (anonymous):

Find the 7th term of the geometric sequence with the given terms: a2 = 768 and a4 = 48; a7 = _____.

OpenStudy (cornzyblack):

u know d formula ryt ? plz write it

OpenStudy (anonymous):

I have it written down, but I'm homeschooled. I learn by watching videos of this guy that explains it. Math is really hard for me and I just don't get it. Today is my last day to finish this class or else I'll fail. I desperately need help. That's why I keep asking. I only have a couple hours left to finish.

OpenStudy (cornzyblack):

ok

OpenStudy (cornzyblack):

an = ar^(n-1) So we have for a2 768 = ar^(2-1) ar^1 = 768 ------ eqn 1 For a4 ar^ (4 -1) = 48 ar^3 = 48 --------- eqn 2 Sp from equation 1, a = 768/ r Substitute (768/r) (r^3) = 48 768r^2 = 48 r^2 = 48/768 r^2 = 1/16 r = 1/4 Now we want to find what a is actually, so using eqn 1 ar^1 = 768 Substituting for r a (1/4) = 768 a = 768 * 4 a = 3072 now for a7 an = ar^(n-1) if a = 3072, n = 7, r = 1/4 a7 = 3072 * (1/4)^ (7-1) a7 =0.75 Am sorrry am kinda slow not that fast at typing

OpenStudy (anonymous):

It is totally fine. You have no idea how much you are helping me

OpenStudy (cornzyblack):

you gat anymore? kuz i will soon be leaving?

OpenStudy (cornzyblack):

u gat anymore?? i will soon be leaving

OpenStudy (anonymous):

Find the indicated sum for the geometric series. S6 for 2 +1 + 1/2 + 1/4 + ...